trentin:

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day

trentin:

Simplify 18−15 = 3

dude:

you first want to get the absolute value by itself So you subtract 15 on both sides \[3\left|x-8\right|=3\] Then you divide by 3 on both sides \[\left|x-8\right|=1\]

dude:

Once again \[x-8=-1\quad \quad \mathrm{or}\quad \:\quad \:x-8=1\]

trentin:

x-8=1

dude:

For the left equation, also add 8 on both sides \[x−8=-1\] Which give you \[x=7\]

dude:

For the equation on the right you also add 8 on both sides \[x−8=1\] which you get \[x=9\]

trentin:

I got the same thing

dude:

Those are your answers! max would be 9 and min would be 7